Q3

The EDTA4- ligand is multidentate with the structural formula shown below. Four of the atoms which can bond to a metal ion are marked with crosses.

Questions:

a) using crosses, indicate where two other atoms could bond in metal ions when EDTA4- forms complexes( 2 marks)

b) explain the term multidentate ligand? (2 marks)

c) The EDTA4- ion reacts with metal(II) ions according to the equation shown below

(M(H2O)6)2+ + EDTA4- —> M(EDTA)2- + 6H2O


in a titration, using a suitable indicator, the following measurements were made:

A solution containing 0.0624 g of hydrated copper(II) sulphate reacted exactly with 25.00 cm3 of 0.0100 M EDTA 4- solution.

using the information, calculate the value of n in the formula of the hydrated copper(II) sulphate, CuSO4 . nH2O (6 marks)

Please top answers 🙏🏾


1
Expert's answer
2022-01-17T10:13:01-0500

(a)






(b)


Multidentate are Ligands with more than one bonded atom.


(c)


Moles of EDTA"^{4-}" used


"=\frac{25}{1000}×0.01=0.00025"


Reacting ratio "=1:1"

"\implies" moles of hydrated copper (ii) Sulphate

"=0.00025"


"0.00025" moles "\equiv0.0624g"

"1\>mole=\frac{0.0624}{0.0025}=249.6"



Molar mass of Anhydrous copper (ii) Sulphate

"=159.6"


Mass of water "=249.6-159.6"

"=90"


Molar mass of water"=18g"

"n=\frac{90}{18}=5"